OpenStudy (1dfanatic):

Solve the system using elimination. Make sure to show/explain your work for credit. 2x + y = 4 2x + 3y = 0

OpenStudy (1dfanatic):

y=4 and 3y=0

OpenStudy (1dfanatic):

idk..

OpenStudy (anonymous):

Subtract the bottom equation from the first equation so you have an equation in one variable. Solve for Y and substitute back into one of the original equations to get x.

OpenStudy (calculusxy):

it's basically like where you subtract the terms that share the same variable. for example, \(2y - 2y\). so you would need to do the same with \(y\) and \(3y\), including \(4\) and \(0\).

OpenStudy (anonymous):

2x + Y = 4 -2x -3Y = 0 -2Y = 4 Solve for Y. Then substitute into 2x + Y = 4 to get x.

OpenStudy (calculusxy):

\(2y - 2y = ?\) \(y - 3y = ?\) \(4 - 0 = ?\)

OpenStudy (anonymous):

Are you good?